OpenStudy (anonymous):

Given that sin x = 0.406 and that x is in quad 1, find sin (x/2)

OpenStudy (anonymous):

\[ \sin(x) = 2 \sin\left ( \frac x 2 \right) \cos\left ( \frac x 2 \right) \] \[ .406 = 2 \sin\left ( \frac x 2 \right)\sqrt{ 1 - \sin^2\left ( \frac x 2 \right)} \] Solve for \[ \sin\left ( \frac x 2 \right) \] to find \[ \sin\left ( \frac x 2 \right)=.21, \sin\left ( \frac x 2 \right)=.98 \] Since we are in the first quadrant, we should only take \[ \sin\left ( \frac x 2 \right)=.21\]

OpenStudy (anonymous):

what identity makes sin x = 2 sin (x/2) cos (x/2) ? i don't get that?

OpenStudy (anonymous):

\[ \sin ( 2 a) = 2 \sin(a) \cos(b)\] and the above comes from \[ \sin(a+b) = \sin(a) \cos(b) + \sin(b) \cos(a) \]